# PriorMapNet: Enhancing Online Vectorized HD Map Construction with Priors

### v1

@@ 4.4 Abalation Study
Table 4: Abalation study of our modules on nuScenes dataset. “PF”, “PPS” and “DMD” denote PF-Encoder, PPS-Decoder and DMD cross-attention, respectively. $u_{t}$ is the total unstable matching score defined in Section 1.
Figure 6: Qualitative results on nuScenes val set. We compare qualitative results PriorMapNet with MapTRv2 and corresponding GTs. Models are trained for 24 epochs. The green area indicates that our method achieves more accurate results.
Abalation study of decoder query prior. We compare various query priors under different BEV feature scales in Tab. 5. MAI in MGMap provides semantic priors for queries, but the improvement is limited. This limitation arises because semantic priors lack positional information crucial for vectorized map elements. We then integrate MAI with query position, but the mAP decreases as the semantic priors are not well-suited for query position. In contrast, our PPS provides prior position and structure, resulting in significant performance improvements. Compared with directly using clustering results as priors, abstracted priors perform better.
Abalation study of PF-Encoder design. We compare various BEV feature priors and encoders, as shown in Tab. 6. All GKT and BEVFormer methods utilize a single-layer architecture. To ensure a fair comparison, we incorporate auxiliary depth supervision across all experimental setups. LSS without prior is the encoder of MapTRv2. The third row denotes a two-layer BEVFormer encoder to exclude the impact of the number of encoder layers. The results show that LSS serves as a good prior generator. Results also demonstrate the effectiveness of Discriminative Loss.
Abalation study of DMD cross-attention design. We compare different decoupling designs of DMD cross-attention, including parallel and serial connection of two MSDA mechanisms, as shown in Tab. 7. In the serial connection, we compare different performance orders of MSDA along multi-sample and multi-scale dimensions. Experimental results demonstrate that our proposed DMD cross-attention improves efficiency and achieves superior performance.
@@
Table 5: Abalation study of decoder query prior. “MAI” denotes Mask-Activated Instance in MGMap, and “MAI-pos” denotes integrating MAI with query position. “PPS-clst” denotes directly using clustering results as priors.
@@
Table 6: Abalation study of the PF-Encoder design. “Dis. Loss” denotes Discriminative Loss, and “-d” denotes auxiliary depth supervision of image features.
@@
Table 7: Abalation study of the DMD cross-attention design. “Sample-then-Scale” denotes performing MSDA along multi-sample then multi-scale, and “Scale-then-Sample” denotes performing MSDA along reverse order. $t_{c}$ denotes the inference time of cross-attention.

### v2

@@ 4.4 Ablation Study
Table 4: Ablation study of our modules on nuScenes dataset. “PF”, “PPS” and “DMD” denote PF-Encoder, PPS-Decoder and DMD cross-attention, respectively. $u_{t}$ is the total unstable matching score defined in Section 1.
Figure 6: Qualitative results on nuScenes val set. We compare visualization results of PriorMapNet with MapTRv2 and corresponding GTs. Models are trained for 24 epochs. The green area indicates that our method achieves more accurate results.
Ablation study of decoder query prior. We compare various query priors under different BEV feature scales in Tab. 5. MAI in MGMap provides semantic priors for queries, but the improvement is limited. This limitation arises because semantic priors lack positional information crucial for vectorized map elements. We then integrate MAI with query position, but the mAP decreases as the semantic priors are not well-suited for query position. In contrast, our PPS provides prior position and structure, resulting in significant performance improvements. Compared with directly using clustering results as priors, abstracted priors perform better.
Ablation study of PF-Encoder design. We compare various BEV feature priors and encoders, as shown in Tab. 6. All GKT and BEVFormer methods utilize a single-layer architecture. To ensure a fair comparison, we incorporate auxiliary depth supervision across all experimental setups. LSS without prior is the encoder of MapTRv2. The third row denotes a two-layer BEVFormer encoder to exclude the impact of the number of encoder layers. The results show that LSS serves as a good prior generator. Results also demonstrate the effectiveness of Discriminative Loss.
Ablation study of DMD cross-attention design. We compare different decoupling designs of DMD cross-attention, including parallel and serial connection of two MSDA mechanisms, as shown in Tab. 7. In the serial connection, we compare different performance orders of MSDA along multi-sample and multi-scale dimensions. Experimental results demonstrate that our proposed DMD cross-attention improves efficiency and achieves superior performance.
@@
Table 5: Ablation study of decoder query prior. “MAI” denotes Mask-Activated Instance in MGMap, and “MAI-pos” denotes integrating MAI with query position. “PPS-clst” denotes directly using clustering results as priors.
@@
Table 6: Ablation study of the PF-Encoder design. “Dis. Loss” denotes Discriminative Loss, and “-d” denotes auxiliary depth supervision of image features.
@@
Table 7: Ablation study of the DMD cross-attention design. “Sample-then-Scale” denotes performing MSDA along multi-sample then multi-scale, and “Scale-then-Sample” denotes performing MSDA along reverse order. $t_{c}$ denotes the inference time of cross-attention.
@@ 5 Conclusion
Supplementary Material
Figure S1: Comparison of the decoders of previous SOTA methods: (a) MapTRv2, (b) MGMap, (c) HIMap, (d) MapQR, (e) InsMapper and (f) our proposed PriorMapNet. (a), (b) and (d) are based on their open-source code, and (c) is based on the implementation details and equations from the paper. For simplicity, we only show the first layer in the transformer decoder. Since InsMapper currently has no open-source code and the specific implementation process is not detailed in the paper, we only show the general structure of its decoder. To highlight the key distinctions between our PriorMapNet and previous SOTA methods, we mark the randomly initialized learnable query position embeddings and reference points in light red, and mark our prior-aware query position embeddings and reference points in yellow.
In the supplementary material, we provide more details, experiments, and analysis of the proposed PriorMapNet, including:
  • More implementation details of our method;
  • Additional ablation studies on hyperparameters;
  • Difference with previous SOTA methods;
  • More qualitative results and failure cases.
S1 More Implementation Details
Feature extraction and multi-modality fusion. Multi-modality inputs consist of multi-view cameras and LiDAR data. For backbones, we utilize ResNet50 (He et al. 2016) for image features extraction and SECOND (Yan, Mao, and Li 2018) for point cloud features extraction. The input image size is 480$\times$800. The voxel size for point cloud is set to [0.1m, 0.1m, 0.2m], and the LiDAR BEV features are upsampled to align with camera BEV features. For BEV feature fusion, we concatenate camera and LiDAR BEV features along feature channels and use a convolution layer to fuse them.
Encoder and decoder. Following MapTRv2 (Liao et al. 2023), we add an auxiliary depth loss for LSS to generate BEV query priors effectively. The PPS-Decoder contains 6 layers to refine the outputs iteratively. In addition to 50 instance-level queries, there are another 300 one-to-many queries that are used to speed up convergence during training. Priors are not used for one-to-many queries.
Training. We use the AdamW optimizer (Loshchilov and Hutter 2017) with an initial learning rate of $6\times 10^{-4}$, and apply the Cosine Annealing learning rate scheduler with a linear warm-up phase (Loshchilov and Hutter 2016). Except for our Discriminative Loss, other training and loss settings keep in line with our baseline MapTRv2.
S2 Additional Ablation Studies
Abalation study on the number of prior queries. We cluster 50 map elements and abstract the position information of the first $N_{\mathrm{pri}}$ elements, as shown in Fig. 2. To most effectively utilize the prior position and structure, we compare the results of different numbers of prior queries (i.e. $N_{\mathrm{pri}}$), as shown in Tab. S1. Utilizing 9 prior queries achieves the best performance. Too few prior queries cannot fully utilize the prior, while too many prior queries limit the learning ability of learnable queries.
$N_{\mathrm{pri}}$ APdiv APped APbou mAP
0 67.5 61.7 65.8 65.0
5 67.2 62.7 67.8 65.9
9 68.7 63.6 68.0 66.8
10 67.3 63.8 68.8 66.7
20 68.1 63.5 68.0 66.5
50 68.4 63.3 67.5 66.4
Table S1: Ablation study of the number of prior queries. Utilizing 9 prior queries achieves the best performance.
Abalation study on the loss weights. The Discriminative Loss is defined as $L_{\mathrm{dis}}=\lambda_{1}L_{\mathrm{var}}+\lambda_{2}L_{\mathrm{dist}}$. Following LaneNet (Neven et al. 2018), we set $\lambda_{1}=\lambda_{2}=\lambda$. In order to obtain the most effective auxiliary discriminative supervision, we compare the results under different loss weights (i.e. $\lambda$), as shown in Tab. S2. Experimental results show the effectiveness of Discriminative Loss, and the best performance can be achieved when $\lambda=1$.
$\lambda$ APdiv APped APbou mAP
0 64.3 60.9 64.8 63.3
0.5 65.6 61.6 63.7 63.6
1 64.9 61.8 64.9 63.9
2 65.1 60.7 64.6 63.5
3 64.5 60.1 65.0 63.2
Table S2: Ablation study of the weight of Discriminative Loss. We set $\lambda_{1}=\lambda_{2}=\lambda$ and setting $\lambda=1$ achieves the best performance.
Abalation study on the borders of the variance and distance loss. $\delta_{\mathrm{v}}$ and $\delta_{\mathrm{d}}$ are respectively the borders of the variance and distance loss. Following LaneNet (Neven et al. 2018), we set $\delta_{\mathrm{d}}=6\delta_{\mathrm{v}}$. To embed the BEV features effectively, we compare the results under different borders of the distance loss (i.e. $\delta_{\mathrm{d}}$), as shown in Tab. S3. Experiments are conducted under $\lambda=1$. Experimental results show that the best performance can be achieved when $\delta_{\mathrm{d}}=3$.
$\delta_{\mathrm{d}}$ APdiv APped APbou mAP
1 64.8 60.5 64.2 63.2
3 64.9 61.8 64.9 63.9
6 66.0 61.2 63.9 63.7
12 65.4 61.0 64.5 63.6
Table S3: Ablation study of the borders of the variance and distance loss. We set $\delta_{\mathrm{d}}=6\delta_{\mathrm{v}}$ and setting $\delta_{\mathrm{d}}=3$ achieves the best performance.
S3 Difference with Previous SOTA Methods
We highlight the key distinctions between our PriorMapNet and previous SOTA methods, including MapTRv2 (Liao et al. 2023), MGMap (Liu et al. 2024a), HIMap (Zhou et al. 2024), MapQR (Liu et al. 2024b) and InsMapper (Xu, Wong, and Zhao 2023). We compare the decoders in Fig. S1, which are the core of these methods.
Motivation and model design. Our motivation and model design diverge from those of previous SOTA methods. PriorMapNet starts from the issue of unstable matching and utilizes pre-computed clustered priors to initialize reference points. Fitted from the map elements in the dataset, prior reference points lower the learning difficulty and achieve stable matching. In contrast, HIMap, MapQR, and InsMapper ignore query priors and use randomly initialized reference points. Based on MapTR series methods, they explore the correlations between instances and points, which introduce additional modules and computational complexity. MGMap proposes Mask-Actived Instance to learn map instance segmentation results and provides semantic priors for instance queries. However, semantic priors lack positional information, which is essential for vectorized map elements. Our PPS-Decoder provides prior position and structure, resulting in significant performance improvements.
Performance. Our PriorMapNet surpasses previous SOTA methods on nuScenes and Argoverse 2 dataset, as illustrated in Tab. 1 and Tab. 2. PriorMapNet achieves SOTA performance under different modalities on nuScenes dataset and different dimensions on Argoverse 2 dataset, demonstrating the generalizability of our method.
S4 Qualitative Results and Failure Cases
We show more qualitative results and failure cases on nuScenes val set in Fig. S2 and Fig. S3, respectively. For each example, the first column represents multi-view camera images. The second column shows the prediction GT. The third and fourth columns respectively show the predicted results of MapTRv2 and our PriorMapNet. Visualization results illustrate that PriorMapNet achieves improved results than MapTRv2. However, in some complex scenes such as intersections, our method fails to predict some map elements, showing our limitations and requiring future work.
Figure S2: Visualization of MapTRv2 and PriorMapNet results and the corresponding GTs. Models are trained for 24 epochs.
Figure S3: Visualization of failure cases. The green areas indicate locations where our method fails to predict accurate results.